d^2 / dx^2 - 2dy / dx= 3e^x sinx by method of undetermined coefficients
1
Expert's answer
2021-08-03T07:08:40-0400

"d^2y / dx^2 - 2dy / dx= 3e^x\sin x"

Find the eigenvectors of the differential operator "d^2\/dx^2 - 2d\/dx". Since this operator is linear and with constant coefficients, the eigenvectors may be found of the following form: "e^{\lambda x}".

"(d^2/dx^2 - 2d/dx)e^{\lambda x}=(\lambda^2-2\lambda)e^{\lambda x}"

Therefore, the null-space o f this operator is generated by the functions "e^{\lambda x}" with such "\lambda", that "\lambda^2-2\lambda=0", i.e. "\lambda_1=0" and "\lambda_2=2".


The general solution of linear ODE is the sum of a partial solution and an arbitrary function from the null space of the differential operator, that is, the function "ae^{2x}+b" with any a,b.


Since "3e^x\sin x" has neither the form "p_1(x)e^{\lambda_1x}", nor the form "p_2(x)e^{\lambda_2x}" with "p_1(x),\,p_2(x)" polynomials on x, but has the form "e^{\lambda x}\sin (ax)", we can seek a partial solution to be of the form

"y=e^x(a_1\sin x+a_2\cos x)"

Let's differentiate:

"y'=e^x(a_1\sin x+a_2\cos x+a_1\cos x-a_2\sin x)=e^x((a_1-a_2)\sin x+(a_1+a_2)\cos x)"

"y''=e^x(((a_1-a_2)-(a_1+a_2))\sin x+((a_1-a_2)+(a_1+a_2))\cos x)="

"=e^x(-2a_2\sin x+2a_1\cos x)"

Hence

"y''-2y'=e^x(-2a_2\sin x+2a_1\cos x)-2e^x((a_1-a_2)\sin x+(a_1+a_2)\cos x)="

"=e^x(-2a_1\sin x-2a_2\cos x)=3e^x\sin x"

Therefore, "a_1=-3\/2", "a_2=0".


Answer. "y=-\frac{3}{2}e^x\sin x+ae^{2x}+b"